OpenStudy (anonymous):

When solving (f of g)(4) do you solve for g first then f or vise versa? The given is f(x)=2x^2-1 and g(x) =6-1/2x^2 please and thank you (:

OpenStudy (anonymous):

(f of g)(x) is f(g(x)), So you should start by solving g(x)=c and plugging that number into f(c). Understand?

OpenStudy (anonymous):

ok let me try this

OpenStudy (anonymous):

i got 8 for g of x, is that right?

OpenStudy (anonymous):

*for (fog)(4)

OpenStudy (anonymous):

so my final answer 127

OpenStudy (anonymous):

let me know if that's right

OpenStudy (anonymous):

so I got it right, correct?

OpenStudy (anonymous):

cause i did get 31

OpenStudy (anonymous):

oops nevermind i go that answer for (gof)(2)

OpenStudy (anonymous):

hmmm... isn't g(4)=-2 ???

OpenStudy (anonymous):

is this correct: \(\large g(x) = 6 - \frac{1}{2}x^2 \) ??

OpenStudy (anonymous):

yes you are right i just realized i miscalculated

OpenStudy (anonymous):

so now you need to calculate f(-2) ....

OpenStudy (anonymous):

so the final answer must be 7

OpenStudy (anonymous):

that's what i got....

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

good work... :)

OpenStudy (anonymous):

wanna try another one and see of we get the same answer?

OpenStudy (anonymous):

(fof)(1)

OpenStudy (anonymous):

f(f(1)) = 1

OpenStudy (anonymous):

I got 1 too

OpenStudy (anonymous):

yay!!!

OpenStudy (anonymous):

now (gof)(2)

OpenStudy (anonymous):

i got 7 :)

OpenStudy (anonymous):

(gog)(0)= 6 ?

OpenStudy (anonymous):

g(f(2)) = g(7) = 6-(49/2) = -37/2 ...:(

OpenStudy (anonymous):

what? i'm confused lol

OpenStudy (anonymous):

for \(\large g(f(2)) = g(2\cdot \color {red}2^2-1)=g(7)=6-\frac{1}{2}\cdot 7^2=6-\frac{49}{2}=\frac{-37}{2} \)

OpenStudy (anonymous):

\(\large g(g(0))=g(6-\frac{1}{2}\cdot \color {red}0^2)=g(6)=6-\frac{1}{2}\cdot 6^2=6-18=-9 \)

OpenStudy (anonymous):

oops.. 6-18 = -12.... haha...

OpenStudy (anonymous):

see I got stuck at 6-1/2*7^2 cause when i multiplied 49 times 1/2 i got 24.5

OpenStudy (anonymous):

yes... that is correct... 49/2 = 24.5

OpenStudy (anonymous):

so 6 - 24.5 = -18.5 = -37/2

OpenStudy (anonymous):

so the answer could be either -18.5 or -37/2?

OpenStudy (anonymous):

they have the same value.... they're equal... it's like saying 3/2 = 1.5

OpenStudy (anonymous):

oh ok i see :)

OpenStudy (anonymous):

can you help me with 4 more problems or are you tired?

OpenStudy (anonymous):

go ahead... post as a new question...

OpenStudy (anonymous):

separately....

OpenStudy (anonymous):

this ones kinda difficult, well for me at least

OpenStudy (anonymous):

i had to take a screenshot of it

OpenStudy (anonymous):

\(\large f(x)=x^2 \) and \(\large g(x)=x^2+1 \) for a), \(\large f(\color {red}{g(x)})=f(\color {red}{x^2+1})=(\color {red}{x^2+1})^2 \) if you want, you can expand this...

OpenStudy (anonymous):

this will give you a polynomial... the domain of all polynomials is all real numbers.

OpenStudy (anonymous):

now you try part b, g(f(x)) = ???

OpenStudy (anonymous):

g(f(x))=(x^2)^2+1 ?

OpenStudy (anonymous):

yes... :)

OpenStudy (anonymous):

so does this mean the same thing as the previous one, all real numbers?

OpenStudy (anonymous):

yep... g(f(x)) is still a polynomial...

OpenStudy (anonymous):

what makes it a polynomial?

OpenStudy (anonymous):

the exponent in each term when expanded is 0 or a positive integer.

OpenStudy (anonymous):

can you give me an example of a non polynomial equation

OpenStudy (anonymous):

in your last problem, (x^2)^2 + 1 = x^4 + 1.... each term has a positive whole number exponent...

OpenStudy (anonymous):

so if an equation has a negative or a fraction that makes it a non-polynomial?

OpenStudy (anonymous):

ok.... non-polynomial: \(\large 45x^3-3x^2+x+7x^{-1} \)

OpenStudy (anonymous):

if what you mean by fraction is like this: \(\large \frac{7}{x} \) , then yes, because notice that \(\large 7x^{-1}=\frac{7}{x} \)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thank you for the info

OpenStudy (anonymous):

now we have (fof)(x) = (x^2)^2

OpenStudy (anonymous):

yw... what about f(f(x) =

OpenStudy (anonymous):

oops.. you're way ahead of me... :) yes.. that's correct...

OpenStudy (anonymous):

im gonna have to bookmark that page you sent me it's very useful

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

google does wonders!

OpenStudy (anonymous):

so we don't have to worry about non-polynomials here huh since the given is pretty much a polynomial already

OpenStudy (anonymous):

yes... anytime you have functions that are polynomials, their composition will always give you polynomials.... i think that's a theorem in algebra but not sure....

OpenStudy (anonymous):

oh ok gotcha

OpenStudy (anonymous):

ok last one for this problem is gog(x) and i got (x^2+1)^2+1

OpenStudy (anonymous):

yep... correct... if u want, you can expand that....

OpenStudy (anonymous):

how so

OpenStudy (anonymous):

(x^2+1)(x^2+1)+1?

OpenStudy (anonymous):

yes... FOIL it out and add the 1....

OpenStudy (anonymous):

x^4+x^2+2

OpenStudy (anonymous):

yep...

OpenStudy (anonymous):

all polynomials will look like this: \(\large a_nx^n+a_{n-1}x^{n-1}+...+a_2x^2+a_1x^1+a_0 \)

OpenStudy (anonymous):

yes, i've seen that before, i was terrified when i first saw that lol

OpenStudy (anonymous):

where the a's are just coefficients....

OpenStudy (anonymous):

i'm still learning about the subs

OpenStudy (anonymous):

that's where i kinda get stuck

OpenStudy (anonymous):

it just means they could be (and usually is) different numbers in front of the x's....

OpenStudy (anonymous):

okayy, ready for the next problem?

OpenStudy (anonymous):

it's just math notation... don't let that worry you...

OpenStudy (anonymous):

g'head...

OpenStudy (anonymous):

\[x^3-x \le0\]

OpenStudy (anonymous):

find the solution set?

OpenStudy (anonymous):

it says solve for the inequality

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok... first you need to solve the equation: \(\large x^3-x=0 \) solve for x .... wha u got???

OpenStudy (anonymous):

u want a hint?

OpenStudy (anonymous):

\[x^3=x\] ?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

no....

OpenStudy (anonymous):

factor that left side completely....

OpenStudy (anonymous):

\(\large x^3-x=x(x^2-1)=x(x+1)(x-1) \) so the equation is: \(\large x(x+1)(x-1)=0 \) now it's obvious what x is...

OpenStudy (anonymous):

\[x(x-1)(x+1)\]

OpenStudy (anonymous):

yes... what's x = ???

OpenStudy (anonymous):

x is 0, 1 and -1?